Solve the differential equation.

dx/ y²(x-y) = dy/ x²(x-y) = dz/ z(x²+y²)
1
Expert's answer
2019-08-26T09:15:18-0400

Here

"P=y^2(x-y), Q=x^2(x-y),Q=z(x^2+y^2)""{dx \over y^2(x-y)}={dy \over x^2(x-y)}={dz \over z(x^2+y^2)}="

From 1st and 2nd fraction, we get 


"(x^2(x-y))dx=(y^2(x-y))dy"

Dividing by "(x-y)," we get


"x^2dx-y^2dy=0""d(x^3-y^3)=0"

Integrating, we get 


"x^3-y^3=c_1"

If we take

"P_1={1 \over y}, Q_1=-{1 \over x},R_1={1 \over z}"


then


"PP_1+QQ_1+RR_1=""={1 \over y}(y^2(x-y))-{1 \over x}(x^2(x-y))+{1 \over z}(z(x^2+y^2))=""=xy-y^2-x^2+xy+x^2+y^2=0"

Thus for the given system of equations, we have 

"{dx \over y^2(x-y)}={dy \over x^2(x-y)}={dz \over z(x^2+y^2)}="


"={dx/y-dy/x+dz/z \over 0}""{dx \over y}=-{dy \over x}={dz \over z}"

From 1st and 2nd fraction, we get

"{dx \over y}=-{dy \over x}""xdx+ydy=0""x^2+y^2=c_2"


Let F be an arbitrary differentiable function. Then the general equation of a partial differential equation is


"F(x^3-y^3,x^2+y^2)=0"